OpenStudy (anonymous):

Factor the polynomial -x^2-9x+22. Plaese help

OpenStudy (anonymous):

-x^2 -11x+2x +22 =-x(x+11) +2(x+11) =(x+11(-x+2) =(x+11)(2-x)